The resonant frequency f of a closed air column is known to depend on the pressure p and the density, of the air and the length, L of the air column. How are these quantities related

1
Expert's answer
2022-02-23T12:11:42-0500

Answer

The relation for resonance frequency is given by

"f=\frac{1}{4L}\sqrt{\frac{K}{\rho}}"

Where K is a coefficient of stiffness

"\rho" is Air density